# Stable generative modeling using Schrödinger bridges

### v2

@@ Abstract.
We consider the problem of sampling from an unknown distribution for which only a sufficiently large number of training samples are available. Such settings have recently drawn considerable interest in the context of generative modelling and Bayesian inference. In this paper, we propose a generative model combining Schrödinger bridges and Langevin dynamics. Schrödinger bridges over an appropriate reversible reference process are used to approximate the conditional transition probability from the available training samples, which is then implemented in a discrete-time reversible Langevin sampler to generate new samples. By setting the kernel bandwidth in the reference process to match the time step size used in the unadjusted Langevin algorithm, our method effectively circumvents any stability issues typically associated with the time-stepping of stiff stochastic differential equations. Moreover, we introduce a novel split-step scheme, ensuring that the generated samples remain within the convex hull of the training samples. Our framework can be naturally extended to generate conditional samples and to Bayesian inference problems. We demonstrate the performance of our proposed scheme through experiments on synthetic datasets with increasing dimensions and on a stochastic subgrid-scale parametrization conditional sampling problem.
Keywords: Langevin dynamics, Schrödinger bridges, generative modeling, Bayesian inference
AMS: 60H10,62F15,62F30,65C05,65C40
@@ 1. Introduction
In this paper, we introduce a new nonparametric approach to generative modeling that combines ideas from Schrödinger bridges [41, 6] and reversible Langevin dynamics [40]. Suppose that we are given $M$ training samples $x^{(i)}\sim\pi$, $i=1,\ldots,M$, from an unknown distribution $\pi$ on $\mathbb{R}^{d}$. Perhaps the simplest nonparametric approach to generative modeling is to build a kernel density estimate (KDE) and then sample from it; the KDE is essentially a mixture model with $M$ components. Alternatively, one could estimate the score function, $s(x)\approx\nabla\log\pi(x)$, without directly estimating $\pi$, and use this estimate as the drift term of Langevin dynamics,
@@
where $W_{\tau}$ denotes standard $d$-dimensional Brownian motion. There is a plethora of ways of estimating the score function [24, 51], and given an estimate for it, one needs to discretize (1), for example using Euler–Maruyama, to obtain an implementable scheme. However, the step size needs to be carefully chosen: a small step size leads to slow convergence, while too large a step yields instability of the numerical scheme, especially for data that are supported on a compact manifold, e.g.,
@@
for some unknown function $g(x)$. Any estimated score function $\hat{s}_{M}(x)$ will take large values for $x$ with $\|g(x)\|^{2}\gg 0$, rendering the Langevin dynamics (1) stiff. This implies that an explicit time integration method such as Euler–Maruyama will require very small step sizes $\Delta\tau>0$.
@@
In this paper, we choose an alternative approach. Instead of first estimating the score function $\hat{s}_{M}(x)$ and then discretising (1) in time, we employ Schrödinger bridges [41, 6] to directly estimate the conditional expectation value
(3) $\mu_{\epsilon}(x):=\mathbb{E}[X_{\epsilon}|X_{0}=x]$
from the given samples $\{x^{(i)}\}_{i=1}^{M}$ for given parameter $\epsilon>0$. We denote this approximation by $m_{\epsilon}(x):\mathbb{R}^{d}\to\mathbb{R}^{d}$. The second key ingredient of our method is to interpret $\epsilon$ as a step size and to read off a Gaussian transition kernel from the Schrödinger bridge approximation in the form of
(4) $X_{n+1}=m_{\epsilon}(X_{n})+\sqrt{\Sigma(X_{n})}\Xi_{n},$
with appropriately defined diffusion matrix $\Sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ and $\Xi_{n}\sim{\rm N}(0,\epsilon I)$. Broadly, $m_{\epsilon}(X_{n})$ controls the drift, while $\sqrt{\Sigma(X_{n})}\Xi_{n}$ introduces noise. An obvious choice for $\Sigma(x)$ is $\Sigma(x)=2I$, which corresponds to (1) and its Euler–Maruyama discretisation with step-size $\Delta\tau=\epsilon$.
In the recursion (4), the step size $\epsilon$ is linked to the error of the Schrödinger bridges approximation. Comparing to directly discretizing (1) using Euler–Maruyama with step-size $\Delta\tau=\epsilon$, we will demonstrate that the scheme (4) is stable and ergodic for all step-sizes $\epsilon>0$ and, hence, $\epsilon$ can be chosen solely on accuracy considerations. Furthermore, the inclusion of the position-dependent diffusion matrix $\Sigma(x)$ makes it better suited for sampling from a manifold.
@@
Our construction utilizes a Sinkhorn scheme [34, 54], which solves the discrete Schrödinger bridge problem of optimally coupling the empirical measure of the training samples with itself over an appropriate diffusion process. By solving the Schrödinger bridge problem, we construct a transition matrix whose state space encompasses all the training points. We generalize this approach to the continuous state space that extends beyond the current training data points. Armed with this transition kernel, which directly leads to the desired approximation $m_{\epsilon}(x)$, we obtain a Markov chain that samples from the underlying distribution of the training data via (4).
@@
From here, we introduce a novel split-step time-stepping scheme, which ensures that the generated samples consistently lie within the convex hull of the training samples and is ergodic [40]. In contrast, using a direct discretization of (1) with Euler–Maruyama results in generating samples on unbounded domains.
@@
for given bandwidth $\rho(x)>0$, which requires appropriate modifications to the Schrödinger bridges considered in [54]. As we demonstrate in our numerical experiments, the resulting sampling scheme (4) provides a better representation of the underlying target distribution. More precisely, we assess the quality of the generated samples using a variable bandwidth kernel and using a fixed bandwidth kernel on synthetic data sets drawn from non-uniform distributions supported on irregular domains and on low-dimensional manifolds. In addition, we explore data-informed choices of $\Sigma(x)$ in (4).
We then extend our method to cover Bayesian inference problems with $\pi$ as prior and to create a conditional generative model. These extensions allow us to perform Bayesian inference in the “simulation-based” setting, i.e., without explicit evaluation of a prior density and, in the case of conditional sampling, even without evaluations of the data likelihood. We demonstrate the performance of our conditional generative model for a stochastic subgrid-scale parametrization problem.
We close this introduction by pointing to a close connection of the considered Schrödinger bridge problems to diffusion maps [8, 7, 39] which have also been employed to approximate the semigroup $\exp(\epsilon\mathcal{L})$, $\epsilon>0$, of a diffusion process with generator $\mathcal{L}$ and invariant distribution $\pi$, from given samples $\{x^{(i)}\}_{i=1}^{M}$. We recall that $\mu_{\epsilon}(x)=\exp(\epsilon\mathcal{L}){\rm Id}(x)$, where ${\rm Id}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the identity map. However, as demonstrated in [54], the Schrödinger bridge approach leads to a more accurate approximation $m_{\epsilon}(x)$ of the analytic $\mu_{\epsilon}(x)$ in terms of $\epsilon>0$.
@@ 1.1. Related work
Langevin dynamics (1) characterizes the motion of particles as they experience a blend of deterministic and stochastic forces. Unlike in this paper, it is typically assumed that the deterministic forcing term $\nabla\log\pi(x)$ is given. Langevin dynamics has been used as a popular tool for sampling data from the target distribution $\pi$. One variation of this is to introduce a symmetric preconditioning operator to the Langevin dynamics and to consider reversible processes of the general form
@@
where $K(x)$ is a symmetric positive definite matrix and we rely on the Itô interpretation of the multiplicative noise term [40]. While (6) has originally been discussed in molecular physics [12, 11, 23], popular choices of $K(x)$ arising from computational statistics include the empirical covariance [13] and the Riemannian metric [14, 30], making this method converge faster and more geometry-aware, while leaving the stationary distribution unchanged. The scaled diffusion matrix (5) corresponds to the choice $K(x)=\rho(x)I$. Optimal choice of $K(x)$ in terms of convergence to equilibrium have recently been discussed in [29].
@@
Our approximation of the semigroup $\exp(\epsilon\mathcal{L})$ relies on recent work on diffusion maps [34, 54], which relies on an accelerated Sinkhorn algorithm. The Sinkhorn algorithm solves for the Markov transition kernel associated with a discrete Schrödinger bridge problem, where the coupling is between the empirical measure of the training samples with itself. This approach results in a symmetric stochastic operator that, notably, approximates the semigroup $\exp(\epsilon\mathcal{L})$ to higher accuracy in $\epsilon$ than standard diffusion map approximations [8, 7]. Separately, the idea of using variable bandwidth kernels can be found very early in the statistics community, for example, in the context of kernel density estimation [46, 52]. Recently, [1] replaces the original fixed bandwidth kernel with the variable bandwidth kernel in the construction of diffusion maps, making the approximation of the generator accurate on unbounded domains. Inspired by this concept, we replace the fixed bandwidth kernel $\Sigma=2I$ with a variable bandwidth kernel (5) for given $\rho(x)>0$. The resulting Schrödinger bridge approximates the semigroup of the reversible Langevin diffusion process (6) with $K(x)=\rho(x)I$.
In recent years, there has been a surge of research interest in the realm of generative modeling. Despite the remarkable achievements of well-established neural network-based generative models, such as variational auto-encoders (VAE) [26, 44], generative adversarial networks (GAN) [15], and diffusion models or score generative models (SGM) [21, 49], they often require meticulous hyperparameter tuning [45, 50] and exhibit a long training time [15, 53]. Furthermore, SGMs solve both a forward and a reverse stochastic differential equation (SDE). The forward SDE introduces noise to the sample, transforming the data into the standard normal distribution, while the reverse SDE takes sample from the standard normal distribution back to the original data distribution, yielding a different sample than the one initially fed into the forward SDE. During the training process, the score function is learned, not for the target distribution, but for the data distribution at each time. Our work, on the other hand, solves only one (forward) SDE, and we learn the score of the target distribution only once.
@@
Several recent studies have combined a range of score function estimation techniques with Langevin dynamics. For example, [49] introduces a noise conditional score network to learn the score function and then uses annealed Langevin dynamics to generate samples. [4] studies the convergence rate of a Langevin based generative model, where the score is estimated using denoising auto-encoders. Such techniques are also studied within the Bayesian imaging community, commonly referred to as “plug and play” [28]. These approaches use neural networks for the estimation of the score function, necessitating substantial fine-tuning, and their effectiveness depends on factors such as the complexity of the approximation families and the architectural structures of the neural networks. In addition, [28] uses an explicit projection to ensure that samples stay on the compact manifold given by (2) which is assumed to be explicitly known. In contrast, we do not assume any knowledge of $\mathcal{M}$.
@@ 1.2. Outline
In Section 2 we construct a Markov chain using a Schrödinger bridge approximation that samples from the given discrete data distribution. In Section 3, we extend this Markov chain to the continuous state space setting by constructing a Gaussian transition kernel which extracts its conditional mean and covariance matrix from the underlying diffusion map approximation. We introduce two discrete-time Langevin samplers; one with an arbitrary data-unaware diffusion and one with a data-aware diffusion matrix in Section 3.1. Theoretical properties such as stability and ergodicity are discussed in Section 3.2. We further discuss the application of variable bandwidth kernels when constructing the Schrödinger bridge in Section 3.3. While Section 3 focuses on finite step-size and finite sample-size implementations, Section 4 establishes connections to the underlying semi-groups and generators in the infinite sample-size limit. We explore the extension of our proposed scheme to a conditional sampling setting and Bayesian inference in Section 5. We demonstrate our proposed methods in Section 6 in a suite of examples, including a conditional sampling exercise with an application to stochastic subgrid-scale parametrization. We conclude in Section 7 with a summary and an outlook.
@@ 2. Discrete Schrödinger bridges
$\mu_{\rm em}({\rm d}x)=\frac{1}{M}\sum_{i=1}^{M}\delta_{x^{(i)}}({\rm d}x)$
in $\mathbb{R}^{d}$ invariant. Here $\delta_{x}({\rm d}x)$ denotes the Dirac delta distribution centred at $x$. In the subsequent section, we will generalise the finding from this section to approximately sample from $\pi$, allowing for the generation of new samples which are different from the given training samples.
We consider the Schrödinger bridge problem of coupling $\mu_{\rm em}$ with itself along a reversible reference process with (unnormalized) transition probabilities
@@
which we collect into a symmetric matrix $T_{\epsilon}\in\mathbb{R}^{M\times M}$. Here $\epsilon>0$ is a tuneable parameter and $K(x)$ is a symmetric positive definite matrix for all $x\in\mathbb{R}^{d}$. Popular choices include $K=I$, $K=\Sigma_{M}$, where $\Sigma_{M}$ is the empirical covariance matrix of the samples $\{x^{(i)}\}_{i=1}^{M}$, and $K=\rho(x)I$, where $\rho(x)>0$ is a scaling function representing variable bandwidth.
Instead of working with the empirical measure $\mu_{\rm em}({\rm d}x)$, we introduce the probability vector $p^{\ast}=(1/M,\ldots,1/M)^{\top}\in\mathbb{R}^{M}$ over $\{x^{(i)}\}_{i=1}^{M}$. Then the associated Schrödinger bridge problem can be reformulated into finding the non-negative scaling vector $v_{\epsilon}\in\mathbb{R}^{M}$ such that the symmetric matrix
(8) $P_{\epsilon}=D(v_{\epsilon})T_{\epsilon}D(v_{\epsilon})$
@@
$P_{\epsilon}p^{\ast}=p^{\ast}.$
Here $D(v)\in\mathbb{R}^{M\times M}$ denotes the diagonal matrix with diagonal entries provided by $v\in\mathbb{R}^{M}$. We remark that the standard scaling used in Schrödinger bridges would lead to a bistochastic matrix $\tilde{P}_{\epsilon}$, which is related to (8) by $\tilde{P}_{\epsilon}=M^{-1}P_{\epsilon}$.
Given $P_{\epsilon}$, one can now construct a Monte Carlo scheme that samples from $\mu_{\rm em}$. Assume the Markov chain is currently in state $x^{(j)}$, then the transition probabilities to the next state $x\in\{x^{(i)}\}_{i=1}^{M}$ are given by
$p_{j}=P_{\epsilon}e_{j}\in\mathbb{R}^{M},$
where $e_{j}\in\mathbb{R}^{M}$ denotes the $j$-th unit vector in $\mathbb{R}^{M}$. Since all entries in $P_{\epsilon}$ are bounded from below provided all samples satisfy $x^{(i)}\in\mathcal{M}$, where $\mathcal{M}$ is a compact submanifold in $\mathbb{R}^{d}$, the constructed Monte Carlo scheme possesses a unique invariant measure given by $p^{\ast}$ and is geometrically ergodic. The rate of convergence can be determined by the diffusion distance
$d(x^{(i)},x^{(j)})=\|p_{i}-p_{j}\|^{2}.$
@@
However, the goal is to approximately sample from the underlying distribution $\pi$ and not just the empirical distribution $\mu_{\rm em}({\rm d}x)$. The required extension of our baseline algorithm is discussed in the following section.

### v3

@@ Abstract.
We consider the problem of sampling from an unknown distribution for which only a sufficiently large number of training samples are available. Such settings have recently drawn considerable interest in the context of generative modelling and Bayesian inference. In this paper, we propose a generative model combining Schrödinger bridges and Langevin dynamics. Schrödinger bridges over an appropriate reversible reference process are used to approximate the conditional transition probability from the available training samples, which is then implemented in a discrete-time reversible Langevin sampler to generate new samples. By setting the kernel bandwidth in the reference process to match the time step size used in the unadjusted Langevin algorithm, our method effectively circumvents any stability issues typically associated with the time-stepping of stiff stochastic differential equations. Moreover, we introduce a novel split-step scheme, ensuring that the generated samples remain within the convex hull of the training samples. Our framework can be naturally extended to generate conditional samples and to Bayesian inference problems. We demonstrate the performance of our proposed scheme through experiments on synthetic datasets with increasing dimensions and on a stochastic subgrid-scale parametrization conditional sampling problem as well as generating sample trajectories of a dynamical system using conditional sampling.
@@ 1. Introduction
In this paper, we introduce a new nonparametric approach to generative modeling that combines ideas from Schrödinger bridges [43, 7] and reversible Langevin dynamics [41]. Suppose that we are given $M$ training samples $x^{(i)}\sim\pi$, $i=1,\ldots,M$, from an unknown distribution $\pi$ on $\mathbb{R}^{d}$. Perhaps the simplest nonparametric approach to generative modeling is to build a kernel density estimate (KDE) and then sample from it; the KDE is essentially a mixture model with $M$ components. Alternatively, one could estimate the score function $s(x)=\nabla\log\pi(x)$, without directly estimating $\pi(x)$, and use this estimate in the Langevin dynamics
@@
where $W_{\tau}$ denotes standard $d$-dimensional Brownian motion [41]. There is a plethora of ways of estimating the score function [26, 53], and given an estimate for it, one needs to discretize (1), for example using Euler–Maruyama, to obtain an implementable scheme. However, the step size needs to be carefully chosen: a small step size leads to slow convergence, while too large a step yields instability of the numerical scheme, especially for data that are supported on or are strongly concentrated about a compact sub-manifold of $\mathbb{R}^{d}$, e.g.,
@@
for some unknown function $g(x)$; a situation commonly encountered in high-dimensional data and referred to as the manifold hypothesis [14, 57] and which is known to be challenging for generative models based on (1) [52]. Any estimated score function, denoted here by $\hat{s}_{M}(x)$, will take large values for $x$ with $\|g(x)\|^{2}\gg 0$, rendering the Langevin dynamics (1) stiff. This implies that an explicit time integration method such as Euler–Maruyama will require very small step sizes $\Delta\tau>0$.
Example 1.1.
In order to illustrate this point consider the singular Gaussian distribution on $\mathbb{R}^{2}$ with $x=(x_{1},x_{2})\in\mathbb{R}^{2}$ satisfying $x_{1}\sim{\rm N}(0,1)$ and $x_{2}=0$. Let us assume that the data samples $x^{(i)}$, $i=1,\ldots,M$, have been polluted by noise from a Gaussian in $\mathbb{R}^{2}$ with mean zero and covariance $\nu I$ for $0<\nu\ll 1$. A standard score estimator $\hat{s}_{M}(x)$ will lead to
(3) $\hat{s}_{\infty}(x)=\left(\begin{array}[]{c}-\frac{1}{1+\nu}x_{1}\\ -\frac{1}{\nu}x_{2}\end{array}\right)$
in the limit $M\to\infty$ and an Euler–Maruyama discretization of the corresponding Langevin dynamics (1) will require step-sizes $\Delta\tau<2\nu$, which become arbitrarily small as $\nu\to 0$. We will return to this simple problem in Example 3.1 where it will be used to demonstrate certain advantages of the methodology proposed in this paper and, in particular, that we can circumvent the just described computational bottleneck as $\nu\to 0$.
@@
Motivated by this illustrative example, we follow an alternative approach in this paper. Namely, instead of first estimating the score function $s(x)$ and then discretising the corresponding approximation to (1) in time, we employ Schrödinger bridges [43, 7] to directly estimate the conditional expectation value
(4) $\mu(x;\epsilon):=\mathbb{E}[X_{\epsilon}|X_{0}=x]$
from the given samples $\{x^{(i)}\}_{i=1}^{M}$ for given parameter $\epsilon>0$. We denote the data-driven estimator by $m(x;\epsilon):\mathbb{R}^{d}\to\mathbb{R}^{d}$.
@@
More specifically, we first note that $\mu(x;\epsilon)=\exp(\epsilon\mathcal{L})\,{\mbox{id}}(x)$, where ${\mbox{id}}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the identity map and $\mathcal{L}$ the generator of the Langevin dynamics (1) [41]. We then approximate the semi-group $\exp(\epsilon\mathcal{L})$ from the given samples $\{x^{(i)}\}_{i=1}^{M}$ using a Schrödinger bridge approximation [35, 58], which optimally couples the empirical measure of the training samples with itself over an appropriate random walk reference process. By solving the Schrödinger bridge problem, we construct a transition matrix whose state space encompasses all the training points, which in turn defines $m(x;\epsilon)$ for all $x=x^{(i)}$, $i=1,\ldots,M$. We then extend this approximation beyond the training data to all $x\in\mathbb{R}^{d}$, which leads to the desired approximation $m(x;\epsilon)$.
The second key ingredient of our method is to interpret $\epsilon$ as a step size and to read off a Gaussian transition kernel from the Schrödinger bridge approximation in the form of
(5) $X_{n+1}=m(X_{n};\epsilon)+\sqrt{\Sigma(X_{n})}\,\Xi_{n},$
with appropriately defined diffusion matrix $\Sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ and $\Xi_{n}\sim{\rm N}(0,\epsilon I)$. Broadly, $m(X_{n};\epsilon)$ controls the drift, while $\sqrt{\Sigma(X_{n})}\,\Xi_{n}$ introduces noise. An obvious choice for $\Sigma(x)$ is $\Sigma(x)=2I$, which corresponds to (1) and its Euler–Maruyama discretisation with step-size $\Delta\tau=\epsilon$. In addition, we explore data-informed choices of $\Sigma(x)$ in (5).
@@
Comparing to directly discretizing (1) using Euler–Maruyama with step-size $\Delta\tau=\epsilon$, we will demonstrate that the scheme (5) is stable and ergodic for all step-sizes $\epsilon>0$ and, hence, $\epsilon$ can be chosen solely on accuracy considerations. We also introduce a novel split-step time-stepping scheme, which ensures that the generated samples lie in the convex hull of the training samples.
@@
for given bandwidth $\rho(x)>0$, which requires appropriate modifications to the Schrödinger bridges considered in [58]. As we demonstrate in our numerical experiments, the resulting sampling scheme (5) provides a better representation of the underlying target distribution. More precisely, we assess the quality of the generated samples using a variable bandwidth kernel and using a fixed bandwidth kernel on synthetic data sets drawn from non-uniform distributions supported on irregular domains and on low-dimensional manifolds.
We then extend our method to cover Bayesian inference problems with $\pi(x)$ as prior and to create a conditional generative model. These extensions allow us to perform Bayesian inference in the “simulation-based” setting, i.e., without explicit evaluation of a prior density and, in the case of conditional sampling, even without evaluations of the data likelihood. We demonstrate the performance of our conditional generative model for a stochastic subgrid-scale parametrization problem and for the generation of synthetic time series of dynamical systems.
@@ 1.1. Related work
Langevin dynamics (1) characterizes the motion of particles as they experience a blend of deterministic and stochastic forces. Unlike in this paper, it is typically assumed that the deterministic forcing term $\nabla\log\pi(x)$ is given. Langevin dynamics has become a popular tool for sampling data from the target distribution $\pi(x)$. One variation of this is to introduce a symmetric preconditioning operator to the Langevin dynamics and to consider reversible processes of the general form
@@
which samples from the distribution $\pi(x)$ for any symmetric positive definite matrix $K(x)$. We adopt here the Itô interpretation of the multiplicative noise term [41]. While (7) has originally been discussed in molecular physics [15, 13, 25], popular choices of $K(x)$ arising from computational statistics include the empirical covariance [16] and the Riemannian metric [17, 31], making this method converge faster and more geometry-aware, while leaving the stationary distribution unchanged. The scaled diffusion matrix (6) corresponds to the choice $K(x)=\rho(x)I$. Optimal choice of $K(x)$ in terms of convergence to equilibrium have recently been discussed in [30].
@@
Our approximation of the semi-group $\exp(\epsilon\mathcal{L})$ relies on recent work on diffusion maps and an accelerated Sinkhorn algorithm [35, 58]. The Sinkhorn algorithm solves for the Markov transition kernel associated with a discrete Schrödinger bridge problem, where the coupling is between the empirical measure of the training samples with itself. This approach results in a symmetric bi-stochastic matrix that, notably, approximates the semi-group $\exp(\epsilon\mathcal{L})$ to higher accuracy in $\epsilon$ than standard diffusion map approximations [9, 8]. Separately, the idea of using variable bandwidth kernels can be found very early in the statistics community, for example, in the context of kernel density estimation [47, 54]. Recently, [1] replaces the original fixed bandwidth kernel with the variable bandwidth kernel in the construction of diffusion maps, making the approximation of the generator accurate on unbounded domains. Inspired by this concept, we replace the fixed bandwidth kernel $\Sigma=2I$ with a variable bandwidth kernel (6) for given $\rho(x)>0$. The resulting Schrödinger bridge approximates the semi-group of the reversible Langevin diffusion process (7) with $K(x)=\rho(x)I$.
@@
Several recent studies have combined a range of score function estimation techniques with Langevin dynamics. For example, [52] introduces a noise conditional score network to learn the score function and then uses annealed Langevin dynamics to generate samples. [4] studies the convergence rate of a Langevin based generative model, where the score is estimated using denoising auto-encoders. Such techniques are also studied within the Bayesian imaging community, commonly referred to as “plug and play” [29]. These approaches typically use neural networks for the estimation of the score function. The Schrödinger bridge sampler proposed in this paper is instead built upon a direct discrete-time approximation to (1). In addition, [29] uses an explicit projection to ensure that samples stay on the compact manifold given by (2) which is assumed to be explicitly known. In contrast, we do not assume any knowledge of $\mathcal{M}$.
We finally mention diffusion models or score generative models (SGM), which have been successfully used for generative modelling, in particular in image generation [23, 52]. These methods solve both a forward and a reverse stochastic differential equation (SDE). The forward SDE introduces noise to the sample, evolving the prior into the standard normal distribution, while the reverse SDE evolves samples from the standard normal distribution back to the original data distribution, yielding a different sample than the one initially fed into the forward SDE. During the training process, the score function is learned for the nonstationary distribution at each time. We mention that Schrödinger bridges have also been implemented in the context of SGMs in order to exactly couple the target distribution with the standard normal distribution [55, 11, 51, 6, 42]. Here we consider a completely different application of Schrödinger bridges; namely the estimation of conditional expectation values from training data.
@@ 1.2. Outline
In Section 2 we construct a Markov chain using a Schrödinger bridge approximation that samples from the given discrete data distribution. In Section 3, we extend this Markov chain to the continuous state space setting by constructing a Gaussian transition kernel which extracts its conditional mean and covariance matrix from the underlying diffusion map approximation. We introduce two discrete-time Langevin samplers; one with a data-unaware diffusion and one with a data-aware diffusion matrix in Section 3.1. Theoretical properties such as stability and ergodicity are discussed in Section 3.2. We further discuss the application of variable bandwidth kernels when constructing the Schrödinger bridge in Section 3.3. Here the goal is to approximate the tails of the reference distribution $\pi(x)$ better from the available training samples. In terms of practical applications, we explore the extension of our proposed scheme to a conditional sampling setting and Bayesian inference in Section 4. We demonstrate our proposed methods in Section 5 in a suite of examples. We start with a couple of synthetic examples demonstrating among others the benefit of variable bandwidth implementations. In terms of applications, we employ conditional sampling to provide a data-driven stochastic subgrid-scale parametrization for multi-scale systems. Furthermore, we show how our approach can be used to generate realistic synthetic trajectories of a dynamical system, given only a single time series for training. We conclude in Section 6 with a summary and an outlook.
@@ 2. Discrete Schrödinger bridges
(8) $\mu_{\rm em}({\rm d}x)=\frac{1}{M}\sum_{i=1}^{M}\delta_{x^{(i)}}({\rm d}x)$
in $\mathbb{R}^{d}$ invariant. Here $\delta_{x}({\rm d}x)$ denotes the Dirac delta distribution centred at $x$. In the subsequent section, we will generalise the finding from this section to approximately sample from $\pi(x)$, allowing for the generation of new samples which are distinct from the given training samples.
We consider the Schrödinger bridge problem of coupling $\mu_{\rm em}({\rm d}x)$ with itself along a reversible reference process with (unnormalized) transition probabilities
@@
which we collect into a symmetric matrix $T\in\mathbb{R}^{M\times M}$. Here $\epsilon>0$ is a tuneable parameter and $K(x)$ is a symmetric positive definite matrix for all $x\in\mathbb{R}^{d}$. Popular choices include $K=I$, $K=\Sigma_{M}$, where $\Sigma_{M}$ is the empirical covariance matrix of the training samples $\{x^{(i)}\}_{i=1}^{M}$, and $K=\rho(x)I$, where $\rho(x)>0$ is a scaling function representing variable bandwidth.
Instead of working with the empirical measure $\mu_{\rm em}({\rm d}x)$, we introduce the probability vector $p^{\ast}=(1/M,\ldots,1/M)^{\top}\in\mathbb{R}^{M}$ over $\{x^{(i)}\}_{i=1}^{M}$. Then the associated Schrödinger bridge problem can be reformulated into finding the non-negative scaling vector $v\in\mathbb{R}^{M}$ such that the symmetric matrix
(10) $P=D(v)TD(v)$
@@
(11) $Pp^{\ast}=p^{\ast}.$
Here $D(v)\in\mathbb{R}^{M\times M}$ denotes the diagonal matrix with diagonal entries provided by $v\in\mathbb{R}^{M}$. We remark that the standard scaling used in Schrödinger bridges would lead to a bi-stochastic matrix $\tilde{P}$, which is related to (10) by $\tilde{P}=M^{-1}P$.
Given $P$, one can now construct a Monte Carlo scheme that samples from $\mu_{\rm em}$. Assume the Markov chain is currently in state $x^{(j)}$, then the transition probabilities to the next state $x\in\{x^{(i)}\}_{i=1}^{M}$ are given by
(12) $p_{j}=Pe_{j}\in\mathbb{R}^{M},$
where $e_{j}\in\mathbb{R}^{M}$ denotes the $j$-th unit vector in $\mathbb{R}^{M}$. Since all entries in $P$ are bounded from below provided all samples satisfy $x^{(i)}\in\mathcal{C}$, where $\mathcal{C}$ is a compact subset in $\mathbb{R}^{d}$, the constructed Monte Carlo scheme possesses a unique invariant measure given by $p^{\ast}$ and is geometrically ergodic. The rate of convergence can be determined by the diffusion distance
(13) $d(x^{(i)},x^{(j)})=\|p_{i}-p_{j}\|^{2}.$
@@
However, the goal is to approximately sample from the underlying distribution $\pi(x)$ and not just the empirical distribution $\mu_{\rm em}({\rm d}x)$. The required extension of our baseline algorithm is discussed in the following section.